OpenStudy (lexy148):

simplify (9 square root of 3)^2

OpenStudy (jhannybean):

\[\large (9\sqrt{3})^2\] Distribute your square value to both terms, the \(9\) and the \(\sqrt{3}\). \[(9)^2\cdot (\sqrt{3})^2\] Remember, \[(\sqrt{3})^2 = \sqrt{3} \cdot \sqrt{3} = 3\]

OpenStudy (jhannybean):

So what would you get?

OpenStudy (lexy148):

i got 81 square root of 3

OpenStudy (jhannybean):

Did you distribute the square to \(\sqrt{3}\)?

OpenStudy (jhannybean):

Look at my first post once more.

OpenStudy (lexy148):

i dont get it still

OpenStudy (jhannybean):

You got the 81 part correct, but multiply \(\sqrt{3} \times \sqrt{3}\). What do you get when you input this into your calculator?

OpenStudy (lexy148):

i get 3

OpenStudy (jhannybean):

Alright, now multiply 3 x 81. What do you get there?

OpenStudy (lexy148):

i get 243

OpenStudy (jhannybean):

Good job :) There you go.

OpenStudy (lexy148):

what do i do after that

OpenStudy (jhannybean):

You're done :D

OpenStudy (lexy148):

oh okay so would it be just 243 or square root of 243

OpenStudy (jhannybean):

Why do you think it would be square root 243?

OpenStudy (lexy148):

idk just because most of the time the answer is usually square root

OpenStudy (jhannybean):

haha, that's not how you're supposed to go about answering these questions!

OpenStudy (lexy148):

oops my bad ill be a little more carful thank you for the help

OpenStudy (jhannybean):

In my first post, I explained when you are squaring TWO terms within the parenthesis, you are DISTRIBUTING the power of 2 to both terms

OpenStudy (jhannybean):

that is why you have \((9)^2 \cdot (\sqrt{3})^2\) . Do you understand that?

OpenStudy (jhannybean):

And remember, don't forget to medal the person you think helped you the most! :)